How do I find the answer to this problem?
Recovering an investment. The manager at Cream of
the Crop bought a load of watermelons for $200. She
priced the melons so that she would make $1.50 profit on
each melon. When all but 30 had been sold, the manager
had recovered her initial investment. How many did she
buy originally?

Answer by scott8148(6628)

x = number of watermelons

" She priced the melons so that she would make $1.50 profit on each melon" ___ p = (200/x) + 1.50

"When all but 30 had been sold, the manager had recovered her initial investment" ___ (x-30)(p) = 200

200x - 6000 + 1.5x^2 - 45x = 200x ___ x^2 - 30x - 4000 = 0

factoring ___ (x - 80)(x + 50) = 0

x = 80